Find the y intercept, x intercept, the axis of symmetry , and vertex of 4= 1/2x^2+4x-5
Can you please help me out? Thanks so much in advance:)
Also can you please show the steps it would really help me understand better:)

Answer by stanbon(75887)

Find the y intercept, x intercept, the axis of symmetry , and vertex of
y = (1/2)x^2+4x-5
------
(1/2)x^2 + 4x = y + 5
-----
(1/2)(x^2 + 8x + 16) = y + 5 +(1/2)16
------
(1/2)(x+4)^2 = y + 13
-------
Vertex: (-4,-13)
Axis of symmetry: x = -4
---
x-int:
Let y = 0 ; solve for "x":
x = [-4 +- sqrt(16 - 4*(1/2)(-5))]/1
----
x = [-4 +- sqrt(26)]
-----------------
y-intercept
Let x = 0, then y = -5
================
Cheers,
Stan H.